# PAR-AdvGAN: Improving Adversarial Attack Capability with Progressive Auto-Regression AdvGAN

### v1

@@ 4.3 Effectiveness Experiment for RQ2: Transferability and Attack Speed
Table 2: The attack success rates (%) on four undefended models and three adversarial trained models by various transferable adversarial attacks. The adversarial examples are crafted on Inc-v3. The best results are in bold.
@@
Table 3: The attack success rates (%) on four undefended models and three adversarial trained models by various transferable adversarial attacks. The adversarial examples are crafted on Inc-v4. The best results are in bold.
@@ 4.3.3 Experiments on IncRes-v2
Table 5: The attack success rates (%) on four undefended models and three adversarial trained models by various transferable adversarial attacks. The adversarial examples are crafted on ResNet-50. The best results are in bold.
@@ 4.3.4 Experiments on ResNet-50
Table 6: The attack success rates (%) on four undefended models and three adversarial trained models by various transferable adversarial attacks. The adversarial examples are crafted on ViT-B/16. The best results are in bold.
@@ 4.3.7 Attack Speed Analysis
As shown in Table. 7, we evaluated the computational efficiency of PAR-AdvGAN and seven competitive baselines using Inc-v3, Inc-v4, and IncRes-v2 as source models. We use FPS as the metric for measuring attack speed. It can be observed that across Inc-v3, Inc-v4, and IncRes-v2, PAR-AdvGAN exhibits speed improvements of 61, 88.3, and 158.5 times over the slowest-performing PGD algorithm among the competitive baselines. Furthermore, in comparison to the fastest-performing FGSM algorithm among the competitive baselines, PAR-AdvGAN achieves speed enhancements of 1.9, 2.5, and 4.4 times, respectively. We assert that PAR-AdvGAN demonstrates significantly higher attack speed in comparison to traditional gradient-based transferable methods, while simultaneously achieving state-of-the-art transferability performance.
@@ 5 Conclusion
In this paper, we present a novel Progressive Auto-Regression AdvGAN (PAR-AdvGAN) algorithm to boost adversarial attack capability through iterative perturbations. Specifically, to address the perturbation escalation issue in AdvGAN, we first adopt a progressive generator network to incorporate the initial sample $x_{0}$ in the perturbation generation process. An auto-regression iterative method is then proposed to include non-initial sample information in generator training. Furthermore, we constrain the distance between initial samples and subsequent samples. Our extensive experimental results exhibit the superior attack transferability of our method. Moreover, compared with the state-of-the-art gradient-based transferable attacks, our method achieves an accelerated attack efficiency.
Code of Ethics and Ethics Statement
All authors of this paper have read and adhered to the ICLR Code of Ethics, as outlined at https://iclr.cc/public/CodeOfEthics. We ensure that our research follows the principles of integrity, fairness, and respect. The methodologies proposed in this work do not involve human subjects, and no personal data is used in our experiments. The datasets and models employed are open-source, and the outcomes of the experiments are aimed at advancing adversarial attack research without promoting any harmful applications. There are no conflicts of interest or external sponsorship affecting this work. We have ensured compliance with ethical standards and legal regulations throughout the research process.
Reproducibility
In order to ensure the reproducibility of our results, we have provided detailed descriptions of the experimental setup in the main text. Our proposed PAR-AdvGAN method is described in Section 3, and the algorithms used are thoroughly outlined, including pseudo-code (Algorithm 1). We have also included hyperparameter settings and model architectures in Section 4.1.3. Additional details about the datasets, metrics, and baseline methods are available in the supplementary materials. Moreover, we have made our code and trained models available in an anonymous repository at https://anonymous.4open.science/r/PAR-01BF/, ensuring easy access for future research and reproduction of our results.
Appendix A Appendix
In this supplementary material, we first show the perturbation comparison diagram of the original AdvGAN and PAR-AdvGAN in Introduction. Secondly, we introduced the optimization objectives for AdvGAN and detailed proof of the omission in Methodology. Furthermore, we state the specific formulas of Attack Success Rate (ASR) and Frames Per Second (FPS) in Metrics. Moreover, we list the specific data table of the experiments performed on IncRes-v2. Finally, we detail the specific parameters of each algorithm in Effectiveness Experiments for code reproduction.
A.1 Schematic diagram of perturbation comparison
Figure 3: Schematic diagram of perturbation comparison (figure. 1 in the main paper)
A.2 Optimization Objectives for AdvGAN
As previously mentioned, in AdvGAN (Xiao et al., 2018), the adversarial sample $x_{adv}$ is synthesized from the original input $x$ and perturbation $G(x)$. By employing a process of iterative and competitive training between the generator $G$ and the discriminator $D$, AdvGAN can generate high-quality adversarial samples. This training approach enables AdvGAN to enhance the generator’s capability to produce perturbations that effectively deceive the discriminator, leading it to classify these perturbations as real samples.
A.2.1 Optimization Objective of Generator
For optimal attack performance, particularly in deceiving the target model $f$ to classify the sample as the target label $l$, AdvGAN utilises the loss function $L_{adv}$ to estimate the likelihood of successfully misleading $f$. In mathematical terms:
$L_{adv}=E_{x}\left(J_{f}(x+G(x),l)\right)$ (16)
Here, $E_{x}$ denotes the expected value of the input data $x$, in accordance with the distribution $p_{data}$. $J_{f}$ represents the loss function employed in training the target model $f$.
A.2.2 Optimization Objective of Discriminator
The discriminator’s role is to distinguish between adversarial samples and real samples. AdvGAN employs the loss function $L_{GAN}$ to maximize the distinction between manipulated data and the real data in discriminator $D$. The mathematical expression of $L_{GAN}$ is as follows:
$L_{GAN}=E_{x}\left(log\left(1-D(x)\right)\right)+E_{x}\left(logD\left(x+G(x)\right)\right)$ (17)
The term $E_{x}\left(logD\left(x+G(x)\right)\right)$ evaluates discriminator $D$’s capability to accurately identify adversarial samples. Conversely, the term $E_{x}\left(log\left(1-D(x)\right)\right)$ assesses the discriminator $D$’s inability to accurately identify original samples $x$.
A.2.3 Optimization Objective of Perturbations
Drawing on the findings from (Carlini & Wagner, 2017; Liu et al., 2016; Isola et al., 2017), AdvGAN incorporates the $L_{hinge}$ loss function to regulate the magnitude of the perturbations. With the incorporating of $L_{hinge}$, AdvGAN effectively limits the perturbation magnitude, ensuring that the generated adversarial samples remain imperceptible and closely resemble the original samples. $L_{hinge}$ is shown as:
$L_{hinge}=E_{x}\left(max(0,\left\|G(x)\right\|_{2}-c)\right)$ (18)
$\left\|\cdot\right\|_{n}$ denotes the $L_{2}$ norm, and $c$ is a user-specified bound.
A.3 Detailed proof of the omission in Eq. 3
Through the chain rule of gradients, we can split $\bigtriangledown_{\theta}$ into $\frac{\partial L_{adv}}{\partial x+G(x)}\cdot\frac{\partial x+G(x)}{\partial G(x)}\cdot\frac{\partial G(x)}{\partial\theta}$. However, in order to explore the relationship between $\frac{\partial L_{adv}}{\partial x+G(x)}$ and $\frac{\partial G(x)}{\partial\theta}$, $\frac{\partial x+G(x)}{\partial G(x)}$ is redundant.
The reason is that, when calculating the partial derivative of $x+G(x)$ with respect to $G(x)$, $x$ is regarded as a constant and $G(x)$ is a variable. Therefore, changes in $x$ have no direct impact on changes in $G(x)$, and its derivative with respect to $G(x)$ is 0, while the derivative of $G(x)$ with respect to itself is 1. Hence, $\frac{\partial x+G(x)}{\partial G(x)}$ equals 1 and can be omitted. In this way, we have removed the redundant terms and unified the remaining two terms.
A.4 Formulas of ASR and FPS
$ASR=\frac{Number\ of\ misleading\ samples}{Number\ of\ adversarial\ samples}$ (19)
$FPS=\frac{Number\ of\ samples}{Running\ time\ of\ these\ samples}$ (20)
Appendix B Ablation Experiment of Iteration Number
(a) 5 iterations
(b) 10 iterations
(c) 15 iterations
Figure 4: Ablation Experiment on the Number of Iterations
B.1 The specific parameters of each algorithm in the effectiveness experiment
B.1.1 Specific parameters for each algorithm on Inc-v3
The specific perturbation rates for PAR-AdvGAN were 8.2869, 9.2728, and 9.9546, while for AdvGAN they were 8.4143, 9.8761, and 10.2608. FGSM had specific perturbation rates of 8.7899, 9.7433, and 10.6936, BIM had rates of 8.4570, 9.4999, and 9.9611, and PGD had rates of 8.7574, 9.7935, and 10.3477. For DI-FGSM, the specific perturbation rates were 8.5112, 9.5550, and 10.0177, for TI-FGSM they were 8.6022, 9.6454, and 10.0996, for MI-FGSM they were 8.9799, 9.7688, and 10.5535, and for SINI-FGSM they were 8.9885, 9.7707, and 10.5512.
Furthermore, the specific parameter configurations for each algorithm under the setting of Inc-v3 are as follows: For AdvGAN, we set $\lambda$ to 0.1 for all three perturbation rates and eps to 10.0, 14.0, and 18.0 respectively. For PAR-AdvGAN, $\lambda$ is set to 1 for all three perturbation rates, and eps is set to 16.0. The number of iterative generations during training is set to 10, and the performance is evaluated separately for one, two, and three iterations of generation for each perturbation rate. For FGSM, we set eps to 9, 10, and 11. For BIM, eps is set to 17, 19, and 20. For PGD, eps is set to 16, 18, and 19. For DI-FGSM, eps is set to 17, 19, and 20. For TI-FGSM, eps is set to 17, 19, and 20. For MI-FGSM, we set eps to 11, 12, and 13. For SINI-FGSM, we set eps to 11, 12, and 13.
B.1.2 Specific parameters for each algorithm on Inc-v4
The specific perturbation rates for PAR-AdvGAN were 9.5509, 11.0456, and 11.6010, while for AdvGAN they were 9.6403, 11.6715, and 12.1057. FGSM had specific perturbation rates of 9.7433, 11.6408, and 12.5848, BIM had rates of 9.9811, 11.4568, and 11.9144, and PGD had rates of 9.7801, 11.3461, and 11.8804. For DI-FGSM, the specific perturbation rates were 10.0124, 11.4927, and 11.9433, for TI-FGSM they were 9.6131, 11.0770, and 11.9980, for MI-FGSM they were 9.8127, 11.4231, and 12.2215, and for SINI-FGSM they were 9.7882, 11.3879, and 12.1840.
Furthermore, the specific parameter configurations for each algorithm under the setting of Inc-v4 are as follows: For AdvGAN, we set $\lambda$ to 0.1 for all three perturbation rates and eps to 12.0, 18.0, and 14.0 respectively. For PAR-AdvGAN, $\lambda$ is set to 1 for all three perturbation rates, and eps is set to 16.0. The number of iterative generations during training is set to 10, and the performance is evaluated separately for one, two, and three iterations of generation for each perturbation rate. For FGSM, we set eps to 10, 12, and 13. For BIM, eps is set to 20, 23, and 24. For PGD, eps is set to 18, 21, and 22. For DI-FGSM, eps is set to 20, 23, and 24. For TI-FGSM, eps is set to 19, 22, and 24. For MI-FGSM, we set eps to 12, 14, and 15. For SINI-FGSM, we set eps to 12, 14, and 15.
B.1.3 Specific parameters for each algorithm on IncRes-v2
The specific perturbation rates for PAR-AdvGAN were 9.8455, 10.5856, and 10.8302, while for AdvGAN they were 10.6939, 11.6411, and 12.5852. FGSM had specific perturbation rates of 10.6939, 11.6411, and 12.5852. BIM had rates of 10.0046, 11.0145, and 11.4804. PGD had rates of 10.3459, 10.8604, and 11.3555. For DI-FGSM, the specific perturbation rates were 10.0459, 11.0495, and 11.5136. For TI-FGSM, they were 10.1527, 10.5860, and 11.1589. For MI-FGSM, they were 10.5966, 11.4195, and 12.2130. For SINI-FGSM, they were 10.5448, 11.3682, and 12.1566.
Furthermore, the specific parameter configurations for each algorithm under the setting of IncRes-v2 are as follows: For AdvGAN, we set $\lambda$ to 0.1 for all three perturbation rates, and eps to 12.0, 18.0, and 28.0, respectively. For PAR-AdvGAN, $\lambda$ is set to 1 for all three perturbation rates, and eps is set to 20.0. The number of iterative generations during training is set to 10, and the performance is evaluated separately for one, two, and three iterations of generation for each perturbation rate. For FGSM, we set eps to 11, 12, and 13. For BIM, eps is set to 20, 22, and 23. For PGD, eps is set to 19, 20, and 21. For DI-FGSM, eps is set to 20, 22, and 23. For TI-FGSM, eps is set to 20, 21, and 22. For MI-FGSM, we set eps to 13, 14, and 15. For SINI-FGSM, we set eps to 13, 14, and 15.
Appendix C Limitations
We notice a slight difference in the attack success rate of PAR-AdvGAN on Inc-v3ens3 and IncResv2ens compared to the best baseline TI-FGSM in Tab. 3, although for overall mASR our method achieves the best performance. This inconsistency suggests a possible dependency of our approach on model selection. The main reason for the inconsistency is related to whether the method requires further interaction with the model after training the Generator regarding the generation process. Our method PAR-AdvGAN doesn’t require the interaction, while TI-FGSM needs to continuously interact with the model for the generation process to obtain gradient information. It is always be feasible for different models and we will consider reducing the dependence on model selection in the future to generate more transferable adversarial examples.

### v3

@@ 4.3 Effectiveness Experiment for RQ2: Transferability and Attack Speed
Table 2: The attack success rates (%) on four undefended models and three adversarial trained models by various transferable adversarial attacks. The adversarial examples are crafted on Inc-v3 with different perturbations. The best results are in bold.
@@
4.3.1 Experiments on Inc-v3
Table 3: The attack success rates (%) on four undefended models and three adversarial trained models by various transferable adversarial attacks. The adversarial examples are crafted on Inc-v4 with different perturbations. The best results are in bold.
@@ 4.3.3 Experiments on IncRes-v2
Table 5: The attack success rates (%) on four undefended models and three adversarial trained models by various transferable adversarial attacks. The adversarial examples are crafted on ResNet-50 with different perturbations. The best results are in bold.
@@ 4.3.4 Experiments on ResNet-50
Table 6: The attack success rates (%) on four undefended models and three adversarial trained models by various transferable adversarial attacks. The adversarial examples are crafted on ViT-B/16 with different perturbations. The best results are in bold.
@@ 4.3.7 Attack Speed Analysis
As shown in Table. 7, we evaluated the computational efficiency of PAR-AdvGAN and seven competitive baselines using Inc-v3, Inc-v4, and IncRes-v2 as source models. We use FPS as the metric for measuring attack speed, representing the number of images that can be processed by the attack per second. It can be observed that across Inc-v3, Inc-v4, and IncRes-v2, PAR-AdvGAN exhibits speed improvements of 61, 88.3, and 158.5 times over the slowest-performing PGD algorithm among the competitive baselines. Furthermore, in comparison to the fastest-performing FGSM algorithm among the competitive baselines, PAR-AdvGAN achieves speed enhancements of 1.9, 2.5, and 4.4 times, respectively. We assert that PAR-AdvGAN demonstrates significantly higher attack speed in comparison to traditional gradient-based transferable methods, while simultaneously achieving state-of-the-art transferability performance.
@@ 5 Conclusion
In this paper, we present a novel PAR-AdvGAN algorithm to boost adversarial attack capability through iterative perturbations. Specifically, to address the perturbation escalation issue in AdvGAN, we first adopt a progressive generator network to incorporate the initial sample $x_{0}$ in the perturbation generation process. An auto-regression iterative method is then proposed to include non-initial sample information in generator training. Furthermore, we constrain the distance between initial samples and subsequent samples. Our extensive experimental results exhibit the superior attack transferability of our method. Moreover, compared with the state-of-the-art gradient-based transferable attacks, our method achieves an accelerated attack efficiency.